# Efficient wireless charging of a quantum battery

### v1

@@ I Introduction
Quantum battery (QB) is a microscopic quantum mechanical device used to temporarily store and release energy in a controllable manner. The idealized procedure of work extraction from a QB is implemented via a cyclic unitary transformation of it governed by the system dynamics plus some controlling fields in a certain time interval. The maximal amount of extractable work in this process is termed as ergotropy EPL , which is defined as the surplus energy of the initial state $\rho$ of the QB with respect to the full set of passive states having the same eigenvalues as $\rho$. Here, a state $\sigma$ is called passive if no work can be extracted from it by using any such unitary transformation passive1 ; passive2 ; passive3 , while it is called active if the maximal work output is nonzero. Moreover, the product of two independent copies of a passive state $\sigma$ may not always be passive ($\sigma$ is called completely passive if $\sigma^{\otimes n}$ is passive with respect to the sum Hamiltonian $H=\sum_{i=1}^{n}H_{i}$, $\forall n$), so the ergotropy for $n$ collective QB cells may be different from the sum of ergotropies for $n$ copies of a single-cell QB EPL ; entangleu .
@@
Up to date, the QB was extensively studied from two main aspects. The first one, centering around the interplay between the performance of a QB and the various quantum resources, revealed that some of its figures of merit (ergotropy, charging power, battery capacity, etc.) are intimately related to quantum characteristics of the battery state such as entropy, coherence, and entanglement workext1 ; workext2 ; qcqb0 ; qcqb1 ; qcqb2 ; qcqb3 ; qcqb4 ; qcqb5 ; shihl1 ; qcqb6 ; correch ; shihl2 ; luomx ; shihl3 . The second one, on the other hand, focused on the charging process of a QB, e.g., the possible realizations of a QB in the spin-chain systems shihl1 ; spin_qb1 ; spin_qb2 ; spin_qb3 ; spin_qb4 , the Tavis-Cummings model qcqb1 ; TCqb1 , the harmonic oscillators noiseqb0 ; oscillator_qb1 ; oscillator_qb2 , the Dicke model Dickeqb1 ; Dickeqb2 ; Dickeqb3 ; Dickeqb4 , and the three-level systems oscillator_qb2 ; threelevelqb . For these battery models, numerous efforts have also been invested into devising feasible schemes to enhance their charging power chpf0 ; chpf1 ; thermalization ; chpf2 ; chpf3 ; chpf4 ; chpf5 ; chpf6 ; chpf7 , exploring their quantum advantages in the charging efficiency (time of charging, input power, total stored energy, etc.) advan1 ; advan2 ; SYKbattery ; advan3 ; advan4 ; advan5 ; advan6 ; advan7 , and identifying bounds on their extractable work and charging power qcqb2 ; bound1 ; bound2 . Moreover, as any quantum system is inevitably disturbed by its surrounding environment, and the charger and QB cannot be excluded from this process too, the dissipative charging of a QB has also been widely explored qcqb4 ; noiseqb0 ; noiseqb1 ; wireless1 ; wireless2 ; noiseqb2 ; PRAppl ; noiseqb3 ; noiseqb4 ; noiseqb5 ; noiseqb6 ; noiseqb7 ; collective ; noiseqb8 ; squeezqb ; twomode ; anjh ; in particular, some structured reservoirs can be exploited as mediators for transferring energy from the charger to a QB, that is, to realize the wireless charging process wireless1 ; wireless2 ; twomode .
@@
Concerning the charging of a QB, our goal is to realize the fast charging and high charging capacity, and it is appealing to provide a general study on the details of the charging process under various situations. However, while most of the previous works considered the case that the charger energy is maximal and the QB is fully discharged initially, the charger might not always be in its fully excited state and one rarely waits until a battery runs out before charging in realistic situations. In fact, by considering different initial states of a QB driven by a classical force, it was found that an optimal charging is achieved if it starts from the ground state initial . Inspired by this, we consider in this paper the general scenarios in which the charger energy is not maximal and the QB is not fully discharged initially, putting emphasis on the role of the number $n$ of charging units and the amount of energy in the charger, the residual ergotropy in the QB, and the coupling strength of the system (i.e., charger $+$ QB) to a common structured reservoir in controlling the charging time and the ergotropy. We will show that the charging performance mediated by the common reservoir improves with the increasing coupling strength, and intuitively, the multiple charging units and the residual ergotropy in the QB may not always be beneficial for shortening the charging time and enhancing the charged ergotropy.
@@
The rest of the paper is organized as follows. In Sec. II, we recall the concept of ergotropy, then in Sec. III, we present the charger-battery model we considered. Sec. IV is devoted to analyzing the charging of a QB via $n$ ($n\geqslant 1$) charging units. Finally, we summarize our findings in Sec. V.
@@ II Preliminaries
$\mathcal{E}_{i}=\max_{V\in\mathcal{U}_{c}}\mathcal{W}(\rho,V),$ (3)
where $V=\sum_{k}e^{-i\phi_{k}}|\varepsilon_{k}\rangle\langle\varepsilon_{\pi_{k}}|\equiv V_{\pi}$, with $\phi_{k}$ being an irrelevant phase factor and $\{\pi_{k}\}_{k=1,\dots,d}$ is a permutation of the elements of $\{1,\dots,d\}$. After optimizing over all possible $\{\pi_{k}\}_{k=1,\dots,d}$, one can obtain that the optimal $V_{\tilde{\pi}}$ ($\tilde{\pi}$ is the optimal permutation) yields $\sigma_{\rho}=V_{\tilde{\pi}}\rho V_{\tilde{\pi}}^{\dagger}=\sum_{kl}\rho_{\tilde{\pi}_{k}\tilde{\pi}_{l}}|\varepsilon_{k}\rangle\langle\varepsilon_{l}|$ qcqb3 , thus
@@
where $\{\rho_{\tilde{\pi}_{k}\tilde{\pi}_{k}}\}_{k=1,\dots,d}$ is the rearrangement of the populations $\{\rho_{kk}\}_{k=1,\dots,d}$ in decreasing order. Equation (4) indicates that $\mathcal{E}_{i}$ also equals to the ergotropy of the dephased state $\delta_{\rho}=\Delta[\rho]=\mathrm{diag}\{\rho_{11},\ldots,\rho_{dd}\}$, i.e., $\mathcal{E}_{i}(\rho)=\mathcal{E}(\delta_{\rho})$ shihl2 , and the passive state associated with $\delta_{\rho}$ is $\tilde{\delta}_{\rho}=\mathrm{diag}\{\rho_{\tilde{\pi}_{1}\tilde{\pi}_{1}},\dots,\rho_{\tilde{\pi}_{d}\tilde{\pi}_{d}}\}$.
@@ III The charger-battery model
where $\omega_{0}$ is the transition frequency between the ground state $|0\rangle$ and excited state $|1\rangle$ of each qubit, $\omega_{0}\sigma_{i}^{+}\sigma_{i}^{-}$ is actually the free Hamiltonian of the $i$th qubit, with $\sigma_{i}^{+}$ ($\sigma_{i}^{-}$) being the Pauli raising (lowering) operator. Moreover, $b_{k}$ ($b_{k}^{\dagger}$) is the annihilation (creation) operator of the field mode $k$ with frequency $\omega_{k}$, and $g_{k}$ is the coupling strength between each qubit and the field mode $k$ of the reservoir. The function $f(t)$, which equals 1 for $t\in(0,\bar{t}]$ and 0 otherwise, is introduced for controlling the switchable coupling of the $N$ qubits to the reservoir, where $\bar{t}$ is the interaction time needed to charge the QB up to its (not necessary the first) dynamical maximum. When $t>\bar{t}$, the interaction between the qubits and the reservoir is switched off. Hereafter we call $\bar{t}$ the charging time.
@@
where $\Omega$ is proportional to the vacuum Rabi frequency and $\lambda$ denotes the frequency width of the spectrum. For such a model, if $N=2$ (the single-charger case) and the “charger $+$ battery” system is initialized in $v_{01}|10\rangle+v_{02}|01\rangle$ ($|v_{01}|^{2}+|v_{02}|^{2}=1$), the dynamics of the system can be obtained analytically common1 ; common2 , and the associated QB has been analyzed in Ref. wireless1 ; wireless2 . For the initial extended Werner-like states, analytical solutions of the two-qubit system can also be obtained in the Laplace transform space pmode0 .
For a general initial state, it is hard to obtain the exact dynamics of the $N$-qubit system, even for $N=2$. So we resort to the pseudomode approach, which describes the coherent interaction between the considered system and the pseudomodes pmode1 ; pmode2 ; pmode3 . Here, the pseudomodes are auxiliary variables defined from the spectrum of the reservoir. For $J(\omega)$ of Eq. (8), this approach results in the following pseudomode master equation in the interaction picture pmode1 ; pmode2 ; pmode3 ; pmode4 ; pmode5 :
@@
$V=\Omega\sum_{i}\sigma_{i}^{+}a+\mathrm{H.c.},$ (10)
by combining of which with Eq. (9), one can solve numerically the dynamics of $\varrho$ without performing any further approximation. After having $\varrho$ at hand, the density operator $\rho$ for the $N$-qubit system can then be obtained via partial tracing of the pseudomode degrees of freedom, and likewise for the charger state $\rho_{\mathrm{ch}}$ and the battery state $\rho_{\mathrm{ba}}$.
@@ IV Charging the QB with different chargers
For the scenario I, the $n$ charging units are not in their excited states, that is, the charger energy is not maximal initially. So for the case of $n=1$, even if the total energy in the charger is transferred to a QB, it cannot be fully charged. But this is a common situation encountered in practice, and it motivates us to consider the charging of an empty QB via $n$ charging units having non-maximal energy, aimed at exploring whether it can complement the lack of energy in a single-charging unit. Here, by saying a QB is empty, we mean that no work can be extracted from it via any cyclic unitary transformation, that is, the battery state is passive. But this does not necessarily mean it is in the ground state. For the scenario II, there is residual ergotropy in the QB (i.e., $\mathcal{E}\neq 0$ initially), which is also a situation we may encounter, as we rarely wait until a battery runs out before charging. In particular, a traditional battery having residual energy may be charged faster than a fully discharged one. But whether this is indeed the case for a QB? This is our motivation for considering the scenario II.
@@ IV.1 The case of a single charging unit
For the case of a single charging unit (i.e., $n=1$), when $c_{1}$ in Eq. (11) equals to 1, the charging of a QB has been investigated in Ref. wireless1 ; wireless2 . For the scenario I with different initial charger states $|\psi\rangle_{c_{1}}$, we show in the left column of Fig. 2 the dynamics of the ergotropy $\mathcal{E}$ as well as its incoherent component $\mathcal{E}_{i}$ and coherent component $\mathcal{E}_{c}$ (in units of $\omega_{0}$) in the good cavity limit $R=20$. For this charger, its initial energy is $c_{1}\omega_{0}$, and the first dynamical maximum of $\mathcal{E}$ is the maximal ergotropy charged on the QB, which, as shown by the different lines in Fig. 2(a), decreases with a decrease in $c_{1}$. This is understandable as the charger energy also decreases with a decrease in $c_{1}$. Moreover, our calculation shows that the energy is transferred to the QB from the very beginning, but as is shown in Fig. 2(a), the ergotropy $\mathcal{E}$ remains zero for a finite time interval ($\lambda t\lesssim 0.1023$ when $R=20$). This shows that the transferred energy cannot be converted into extractable work immediately in certain situations. As for the incoherent and coherent contributions to the ergotropy, by comparing the left three panels of Fig. 2, one can see that for the charger with a large $c_{1}$, the incoherent contribution to ergotropy is dominant, whereas for that with a small $c_{1}$, the coherent contribution to ergotropy turns to be dominant.
@@
When considering the scenario II, we show in the right column of Fig. 2 the dynamics of $\mathcal{E}$, $\mathcal{E}_{i}$, and $\mathcal{E}_{c}$ (in units of $\omega_{0}$) in the good cavity limit $R=20$. For this scenario, the initial state $|\varphi\rangle_{e_{1}}$ of the QB is active and the residual ergotropy in it is $e_{1}\omega_{0}$. From Fig. 2(b) one can see that $\mathcal{E}$ first decays to zero (i.e., fully discharged) and then turns to be increased to its first dynamical maximum, and this can be recognized as the maximal ergotropy charged on the QB as the subsequent dynamical maxima become smaller and smaller. This phenomenon is in contrast to that of a traditional battery, as it indicates that when there is residual ergotropy in the QB, it will inevitably undergo a fully discharging process, after which it can then be continuously charged. Moreover, for a fixed $R$, the first dynamical maximum of $\mathcal{E}$ may be smaller than its initial value $e_{1}\omega_{0}$ when $e_{1}$ is larger than a critical value (e.g., $e_{1}\gtrsim 0.7129$ when $R=20$), that is, the QB cannot be further charged in this case. Of course, the critical $e_{1}$ can be increased by increasing the coupling strength. By comparing the right three panels of Fig. 2, one can see that for this scenario, the incoherent contribution to ergotropy is always dominant, which is somewhat different from that of the scenario I.
@@
In the above we have analyzed charging of a QB in the good cavity limit. Next, we see the case of the bad cavity limit. As is shown in Fig. 3(a), the ergotropy $\mathcal{E}$ for the scenario I increases very slowly with the increase of $\lambda t$, and the asymptotic value is much less than that in the good cavity limit. Physically, such an inefficiency has its roots in the weak coupling of the qubits to the reservoir, which induces a very weak indirect interaction between the the charger and the QB, so most of the energy in the charger is leaked into the reservoir. In particular, when $c_{1}=1$ (i.e., the charger is in its excited state), the QB cannot be charged at all. But this does not mean that there is no energy being transferred to the QB. In fact, in the infinite time limit one has $E_{\mathrm{ba}}(\infty)=\mathrm{tr}[\rho_{\mathrm{ba}}(\infty)H_{\mathrm{ba}}]=c_{1}\omega_{0}/4$ ($\forall c_{1}$) [it equals the injected energy as $E_{\mathrm{ba}}(0)=0$], where $\rho_{\mathrm{ba}}(t)$ and $H_{\mathrm{ba}}$ are the density operator and free Hamiltonian of the QB, respectively. But when $c_{1}=1$, the energy $E_{\mathrm{ba}}$ cannot be extracted by means of any cyclic unitary transformation. This is because for this case the battery state $\rho_{\mathrm{ba}}=\mathrm{diag}\{e(t),1-e(t)\}$, where $e(t)<0.25$ ($\forall t$). We have also calculated the incoherent and coherent components of $\mathcal{E}$ (for the conciseness of this paper, we do not plot them here) and it is found that $\mathcal{E}\equiv\mathcal{E}_{c}$. As for the scenario II, although the initial battery state $|\varphi\rangle_{e_{1}}$ possesses nonzero ergotropy, from Fig. 3(b) one can note that $\mathcal{E}$ always decays with the evolving time, thereby the QB cannot be charged in this case.
@@
Figure 4: The $R$ dependence of the charging time $\lambda\bar{t}$ and ergotropy $\bar{\mathcal{E}}$ charged by a single charging unit at time $\bar{t}$. The left (right) two panels are plotted with the charger and QB being in the initial states $|\psi\rangle_{c_{1}}$ and $|0\rangle$ ($|1\rangle$ and $|\varphi\rangle_{e_{1}}$), respectively. For small $R$, $\lambda\bar{t}$ is very large and the top two panels are cut to better visual its behavior in the large $R$ region, and the first five points in (c) and (d) are plotted with $R=0.1$, 1, 2, 3, and 4, respectively.
@@
In the following, we examine charging rate of the QB characterized by the charging time $\bar{t}$. As is shown in the top two panels of Fig. 4, for the scenario I, $\lambda\bar{t}$ is independent of $c_{1}\neq 0$ and it decreases monotonically with the increase of $R$. In the bad cavity limit, $\lambda\bar{t}$ will be infinitely large. On the contrary, it is very small in the good cavity limit (e.g., $\lambda\bar{t}\simeq 0.0314$ for $R=100$), that is, a fast charging is achieved. For the scenario II, $\lambda\bar{t}$ is weakly dependent on $e_{1}$, and for a given $e_{1}$ it also decreases monotonically with the increasing $R$. This shows that the charging rate of a QB can be enhanced efficiently by increasing the coupling strength between the qubits (i.e., charger $+$ QB) and the reservoir. Moreover, from Fig. 4(a) and (b) one can note that the charging times for the scenarios I and II are approximately the same, that is, the residual ergotropy in the QB does not help to improve the charging rate.

### v2

@@ I Introduction
Quantum battery (QB) is a microscopic quantum mechanical device used to temporarily store and release energy in a controllable manner. The idealized procedure of work extraction from a QB is implemented via a cyclic unitary transformation of it governed by the system dynamics plus some controlling fields in a certain time interval. The maximal amount of extractable work in this process is termed as ergotropy EPL , which is defined as the surplus energy of the initial state $\rho$ of the QB with respect to the full set of passive states having the same eigenvalues as $\rho$. Here, a state $\sigma$ is called passive if no work can be extracted from it by using any such unitary transformation passive1 ; passive2 ; passive3 , while it is called active otherwise. Moreover, the product of two independent copies of a passive state $\sigma$ may not always be passive ($\sigma$ is called completely passive if $\sigma^{\otimes n}$ is passive with respect to the sum Hamiltonian $H=\sum_{i=1}^{n}H_{i}$, $\forall n$), so the ergotropy for $n$ collective QB cells may be different from the sum of ergotropies for $n$ copies of a single-cell QB EPL ; entangleu .
@@
To date, the QB has been extensively studied from two main aspects. The first one, centering around the interplay between the performance of a QB and the various quantum resources, revealed that some of its figures of merit (ergotropy, charging power, battery capacity, etc.) are intimately related to quantum characteristics of the battery state such as entropy, coherence, and entanglement workext1 ; workext2 ; qcqb0 ; qcqb1 ; qcqb2 ; qcqb3 ; qcqb4 ; qcqb5 ; shihl1 ; qcqb6 ; correch ; shihl2 ; luomx ; shihl3 ; AVS . The second one, on the other hand, focused on the charging process of a QB, e.g., the possible realizations of a QB in the spin-chain systems shihl1 ; spin_qb1 ; spin_qb2 ; spin_qb3 ; spin_qb4 ; spin_qb5 , the Tavis-Cummings model qcqb1 ; TCqb1 , the harmonic oscillators noiseqb0 ; oscillator_qb1 ; oscillator_qb2 , the Dicke model Dickeqb1 ; Dickeqb2 ; Dickeqb3 ; Dickeqb4 , and the three-level systems oscillator_qb2 ; threelevelqb ; Dou1 ; Dou2 . For these battery models, numerous efforts have also been invested into devising feasible schemes to enhance their charging power chpf0 ; chpf1 ; thermalization ; chpf2 ; chpf3 ; chpf4 ; chpf5 ; chpf6 ; chpf7 ; Shangc , exploring their quantum advantages in the charging efficiency (time of charging, input power, total stored energy, etc.) advan1 ; advan2 ; SYKbattery ; advan3 ; advan4 ; advan5 ; advan6 ; advan7 , and identifying bounds on their extractable work and charging power qcqb2 ; bound1 ; bound2 . Moreover, as any quantum system is inevitably disturbed by its surrounding environment, and the charger and QB cannot be excluded from this process too, the dissipative charging of a QB has also been widely explored qcqb4 ; noiseqb0 ; noiseqb1 ; wireless1 ; wireless2 ; noiseqb2 ; PRAppl ; noiseqb3 ; noiseqb4 ; noiseqb5 ; noiseqb6 ; noiseqb7 ; collective ; noiseqb8 ; squeezqb ; twomode ; anjh ; weak ; in particular, some structured reservoirs can be exploited as mediators for transferring energy from the charger to a QB, that is, to realize the wireless charging process wireless1 ; wireless2 ; twomode .
@@
Concerning the charging of a QB, our goal is to realize the fast charging and high charging capacity, and it is appealing to provide a general study on the details of the charging process under various situations. However, while most of the previous works considered the case where the charger energy is maximal and the QB is fully discharged initially, the charger might not always be in its fully excited state and one rarely waits until a battery runs out before charging in realistic situations. In fact, by considering different initial states of a QB driven by a classical force, it was found that an optimal charging is achieved if it starts from the ground state initial . Inspired by this, we consider in this paper the general scenarios in which the charger energy is not maximal and the QB is not fully discharged initially, putting emphasis on the role of the number $n$ of charging units and the amount of energy in the charger, the residual ergotropy in the QB, and the coupling strength of the system (i.e., charger $+$ QB) to a common structured reservoir in controlling the charging time and the ergotropy. We will show that the charging performance mediated by the common reservoir improves with the increasing coupling strength, and counter intuitively, the multiple charging units and the residual ergotropy in the QB may not always be beneficial for shortening the charging time and enhancing the charged ergotropy.
@@
The rest of the paper is organized as follows. In Sec. II we recall the concept of ergotropy, and in Sec. III we present the charger-battery model we considered. Section IV is devoted to analyzing the charging of a QB via $n$ ($n\geqslant 1$) charging units. Finally, we summarize our findings in Sec. V.
@@ II Preliminaries
$\mathcal{E}_{i}=\max_{V\in\mathcal{U}_{c}^{(i)}}\mathcal{W}(\rho,V),$ (3)
where $V=\sum_{k}e^{-i\phi_{k}}|\varepsilon_{k}\rangle\langle\varepsilon_{\pi_{k}}|\equiv V_{\pi}$, with $\phi_{k}$ being an irrelevant phase factor and $\{\pi_{k}\}_{k=1,\dots,d}$ being a permutation of the elements of $\{1,\dots,d\}$. After optimizing over all possible $\{\pi_{k}\}_{k=1,\dots,d}$, one can obtain that the optimal $V_{\tilde{\pi}}$ ($\tilde{\pi}$ is the optimal permutation) yields $\sigma_{\rho}=V_{\tilde{\pi}}\rho V_{\tilde{\pi}}^{\dagger}=\sum_{kl}\rho_{\tilde{\pi}_{k}\tilde{\pi}_{l}}|\varepsilon_{k}\rangle\langle\varepsilon_{l}|$ qcqb3 , thus
@@
where $\{\rho_{\tilde{\pi}_{k}\tilde{\pi}_{k}}\}_{k=1,\dots,d}$ is the rearrangement of the populations $\{\rho_{kk}\}_{k=1,\dots,d}$ in decreasing order. Equation (4) indicates that $\mathcal{E}_{i}$ also equals the ergotropy of the dephased state $\delta_{\rho}=\Delta[\rho]=\mathrm{diag}\{\rho_{11},\ldots,\rho_{dd}\}$, i.e., $\mathcal{E}_{i}(\rho)=\mathcal{E}(\delta_{\rho})$ shihl2 , and the passive state associated with $\delta_{\rho}$ is $\tilde{\delta}_{\rho}=\mathrm{diag}\{\rho_{\tilde{\pi}_{1}\tilde{\pi}_{1}},\dots,\rho_{\tilde{\pi}_{d}\tilde{\pi}_{d}}\}$.
@@ III The charger-battery model
where $\omega_{0}$ is the transition frequency between the ground state $|0\rangle$ and excited state $|1\rangle$ of each qubit and $\omega_{0}\sigma_{i}^{+}\sigma_{i}^{-}$ is actually the free Hamiltonian of the $i$th qubit, with $\sigma_{i}^{+}$ ($\sigma_{i}^{-}$) being the Pauli raising (lowering) operator. Moreover, $b_{k}$ ($b_{k}^{\dagger}$) is the annihilation (creation) operator of the field mode $k$ with frequency $\omega_{k}$, and $g_{k}$ is the coupling strength between each qubit and the field mode $k$ of the reservoir. The function $f(t)$, which equals 1 for $t\in(0,\bar{t}]$ and 0 otherwise, is introduced for controlling the switchable coupling of the $N$ qubits to the reservoir, where $\bar{t}$ is the interaction time needed to charge the QB up to its (not necessarily the first) dynamical maximum. When $t>\bar{t}$, the interaction between the qubits and the reservoir is switched off. Hereafter we call $\bar{t}$ the charging time.
@@
where $\Omega$ is the effective coupling strength proportional to the vacuum Rabi frequency, $\lambda$ denotes the frequency width of the spectrum, and the ideal cavity limit (no losses) is obtained at $\lambda\rightarrow 0$. For this model, if $N=2$ (i.e., the single charger case) and the “charger $+$ battery” system is initialized in $v_{01}|10\rangle+v_{02}|01\rangle$ ($|v_{01}|^{2}+|v_{02}|^{2}=1$), the dynamics of the system can be obtained analytically common1 ; common2 , and the associated QB has been analyzed in Refs. wireless1 ; wireless2 . For the initial extended Werner-like states, analytical solutions of the two-qubit system can also be obtained in the Laplace transform space pmode0 .
For a general initial state, it is hard to obtain the exact dynamics of the $N$-qubit system, even for $N=2$. So, we resort to the pseudomode approach, which describes the coherent interaction between the considered system and the pseudomodes pmode1 ; pmode2 ; pmode3 . Here, the pseudomodes are auxiliary variables defined from the spectrum of the reservoir. For $J(\omega)$ of Eq. (8), this approach results in the following pseudomode master equation in the interaction picture pmode1 ; pmode2 ; pmode3 ; pmode4 ; pmode5 :
@@
$V=\Omega\sum_{i}\sigma_{i}^{+}a+\mathrm{H.c.};$ (10)
by combining this with Eq. (9), one can solve numerically the dynamics of $\varrho$ without performing any further approximation. After having $\varrho$ at hand, the density operator $\rho$ for the $N$-qubit system can then be obtained via partial tracing of the pseudomode degrees of freedom, and likewise for the charger state $\rho_{\mathrm{ch}}$ and the battery state $\rho_{\mathrm{ba}}$.
@@ IV Charging the QB with different chargers
For scenario I, the $n$ charging units are not in their excited states, that is, the charger energy is not maximal initially. So, for the case of $n=1$, even if the total energy in the charger is transferred to a QB, it cannot be fully charged. But, this is a common situation encountered in practice, and it motivates us to consider the charging of an empty QB via $n$ charging units having nonmaximal energy, aimed at exploring whether it can complement the lack of energy in a single charging unit. Here, by saying a QB is empty, we mean that no work can be extracted from it via any cyclic unitary transformation, that is, the battery state is passive. But, this does not necessarily mean it is in the ground state. For scenario II, there is residual ergotropy in the QB (i.e., $\mathcal{E}\neq 0$ initially), which is also a situation we may encounter, as we rarely wait until a battery runs out before charging. In particular, a traditional battery having residual energy may be charged faster than a fully discharged one. But is this indeed the case for a QB? This is our motivation for considering scenario II.
@@ IV.1 The case of a single charging unit
For the case of a single charging unit (i.e., $n=1$), if $c_{1}$ in Eq. (11) is equal to 1, the charging of a QB has been investigated in Refs. wireless1 ; wireless2 . For scenario I with different initial charger states $|\psi\rangle_{c_{1}}$, we show in the left column of Fig. 2 the dynamics of the ergotropy $\mathcal{E}$ as well as its incoherent component $\mathcal{E}_{i}$ and coherent component $\mathcal{E}_{c}$ (in units of $\omega_{0}$) in the good cavity limit $R=20$. For this charger, its initial energy is $c_{1}\omega_{0}$, and the first dynamical maximum of $\mathcal{E}$ corresponds to the maximal ergotropy charged on the QB, which, as shown by the different lines in Fig. 2(a), decreases with a decrease in $c_{1}$. This is understandable as the charger energy also decreases with a decrease in $c_{1}$. Moreover, our calculation shows that the energy is transferred to the QB from the very beginning, but as is shown in Fig. 2(a), the ergotropy $\mathcal{E}$ remains 0 for a finite time interval ($\lambda t\lesssim 0.1023$ when $R=20$). This shows that the transferred energy cannot be converted into extractable work immediately in certain situations. As for the incoherent and coherent contributions to the ergotropy, by comparing the left three panels of Fig. 2, one can see that $\mathcal{E}_{c}\equiv 0$ for the charger with $c_{1}=1$. Moreover, if $10.7073\lesssim c_{1}<$ 1, the incoherent and coherent components of the ergotropy dominate its dynamics alternately, and if $c_{1}\lesssim 0.7073$, the coherent contribution will always be dominant.
@@
When considering scenario II, we show in the right column of Fig. 2 the dynamics of $\mathcal{E}$, $\mathcal{E}_{i}$, and $\mathcal{E}_{c}$ (in units of $\omega_{0}$) in the good cavity limit $R=20$. For this scenario, the initial state $|\varphi\rangle_{e_{1}}$ of the QB is active if $e_{1}\neq 0$ and the residual ergotropy in it is $e_{1}\omega_{0}$. From Fig. 2(b) one can see that $\mathcal{E}$ first decays to 0 (i.e., fully discharged) and then turns to be increased to its first dynamical maximum, and this can be recognized as the maximal ergotropy charged on the QB as the subsequent dynamical maxima become smaller and smaller. This phenomenon is in contrast to that of a traditional battery, as it indicates that when there is residual ergotropy in the QB, it will inevitably undergo a fully discharging process, after which it can then be continuously charged. Moreover, for a fixed $R$, the first dynamical maximum of $\mathcal{E}$ may be smaller than its initial value $e_{1}\omega_{0}$ when $e_{1}$ is larger than a critical value (e.g., $e_{1}\gtrsim 0.7129$ when $R=20$), that is, the QB cannot be further charged in this case. Of course, the critical $e_{1}$ can be increased by increasing the coupling strength. By comparing the right three panels of Fig. 2, one can see that for this scenario, $\mathcal{E}_{c}\equiv 0$ when $e_{1}=0$, and when $e_{1}\in(0,1)$, the incoherent and coherent components of the ergotropy $\mathcal{E}$ dominate its dynamics alternately. Similar to that of scenario I, there is always a moment when $\mathcal{E}$ is zero or infinitesimal. Physically, this phenomenon for both scenarios is due to the fact that the indirect coupling between the charger and the QB is induced by the cavity mode through the exchange of cavity photons, thus there is a competition between the energy flow from the charger to the QB and energy backflow to the charger, both via the mediation of the cavity field. In fact, at the time when $\mathcal{E}\simeq 0$, much energy flows back to the charger and the cavity field. Moreover, as shown in Fig. 2, there is a series of peaks, and as time goes by, the intensity becomes smaller and smaller. This is understandable as with the evolving of time, more and more energy is dissipated into the cavity field, and in the infinite time limit, the dissipated energy is $c_{1}\omega_{0}/2$ for scenario I and $(1+3e_{1})\omega_{0}/2$ for scenario II.
@@
In the above we have analyzed charging of a QB in the good cavity limit. Next, we see the case of the bad cavity limit. As shown in Fig. 3(a), the ergotropy $\mathcal{E}$ for scenario I increases very slowly with the increase of $\lambda t$, and the asymptotic value is much smaller than $\bar{\mathcal{E}}$ in the good cavity limit. Physically, such an inefficiency has its roots in the weak coupling of the qubits to the reservoir, which induces a very weak indirect interaction between the charger and the QB, thus most of the energy in the charger is leaked into the reservoir. In particular, when $c_{1}=1$ (i.e., the charger is in its excited state), the QB cannot be charged at all. But, this does not mean that there is no energy being transferred to the QB, as for this case $\rho_{\mathrm{ba}}=\mathrm{diag}\{|\nu_{2}|^{2},1-|\nu_{2}|^{2}\}$ (see Appendix A), and one can show that $|\nu_{2}|$ increases from 0 to 0.5 as the time evolves, which yields a nonzero mean energy $E_{\mathrm{ba}}=|\nu_{2}|^{2}\omega_{0}$ of the QB and a vanishing ergotropy as $\mathcal{E}=\max\{0,2|\nu_{2}|^{2}-1\}\equiv 0$. Hence, the energy cannot be extracted by means of any cyclic unitary operation. We have also calculated the incoherent and coherent components of $\mathcal{E}$ and it is found that $\mathcal{E}\equiv\mathcal{E}_{c}$. For scenario II, although the initial QB possesses nonzero ergotropy, from Fig. 3(b) one can see that $\mathcal{E}$ always decays with the evolving time, so the QB cannot be charged in this case.
@@
Figure 4: The $R$ dependence of the charging time $\lambda\bar{t}$ and ergotropy $\bar{\mathcal{E}}$ charged by a single charging unit at time $\bar{t}$. The left (right) two panels are plotted with the charger and QB being in the initial states $|\psi\rangle_{c_{1}}$ and $|0\rangle$ ($|1\rangle$ and $|\varphi\rangle_{e_{1}}$), respectively. For small $R$, $\lambda\bar{t}$ is very large and the top two panels are cut to better show its behavior in the large $R$ region, and the first five points in (c) and (d) are plotted with $R=0.1$, 1, 2, 3, and 4, respectively.
@@
In the following, we examine the charging rate of the QB characterized by the charging time $\bar{t}$. As is shown in the top two panels of Fig. 4, for scenario I, $\lambda\bar{t}$ is independent of $c_{1}\neq 0$ and it decreases monotonically with the increase of $R$. In the bad cavity limit, $\lambda\bar{t}$ will be infinitely large. On the contrary, it is very small in the good cavity limit (e.g., $\lambda\bar{t}\simeq 0.0314$ for $R=100$), that is, a fast charging is achieved. In fact, for scenario I, both the ergotropy charged on the QB and the charging time can be obtained analytically (see Appendix A); specifically, $\bar{t}=2\pi/|\zeta|$ with $\zeta=\lambda(1-4R^{2})^{1/2}$. Hence, it is the same for different $c_{1}\neq 0$. For scenario II, $\bar{t}$ is weakly dependent on $e_{1}$, and for a given $e_{1}\neq 1$ it also decreases monotonically with the increase of $R$. This shows that the charging rate of a QB could be efficiently enhanced by increasing the coupling strength between the charger-battery system and the reservoir. Moreover, from Figs. 4(a) and (b) one can note that the charging times for scenarios I and II are approximately the same, that is, the residual ergotropy in the QB does not help to improve the charging rate. For scenario II, however, it is hard to obtain an analytical solution of the ergotropy $\mathcal{E}$. Physically, the insensitivity of the charging time $\bar{t}$ to the initial charger and battery states is due to the fact that the indirect coupling of the charger and QB is induced by the cavity mode through the exchange of cavity photons, and it is the effective coupling strength $\Omega$ rather than the initial system state that dominates the cavity losses (photon escape rate) common1 ; common2 , thus it is understandable that $\bar{t}$ is insensitive to the initial system state. The decrease of $\bar{t}$ with the increasing $R$ can also be explained from this perspective as $R=\sqrt{2}\Omega/\lambda$. The ergotropy charged on the QB strongly depends on the initial charger state $|\psi\rangle_{c_{1}}$, the reason for which is that different charger states have different amounts of internal energy.